# On dissipation operators of Quantum Optics

### v2

@@ Abstract
We consider dissipation operators used in Quantum Optics for the description of quantum spontaneous emission in the context of damped driven Jaynes–Cummings equations. The equations describe quantised one-mode Maxwell field coupled to a two-level molecule. Our main result is the symmetry and nonpositivity of basic dissipation operator of Quantum Optics.
@@
Keywords: Quantum Optics; laser; Jaynes–Cummings model; Hamiltonian; density operator; pumping; dissipation; Hilbert–Schmidt operator; trace.
@@ 1 Jaynes–Cummings equation
The Jaynes–Cummings equation (JCE) is basic model of Quantum Optics. The model without damping and pumping was introduced in 1963 by Jaynes and Cummings [12] (the survey can be found in [16]). The damped driven Jaynes–Cummings equations read as [21, (5.107)]:
@@ 2 Dissipation operators and main results
We will consider the dissipation operators [21, (5.107)] and [5, (5.6.20)]:
@@
$\rm HS$} is the Hilbert space of Hermitian Hilbert–Schmidt operators with the inner product [19, Thm VI.22 (c)]
@@
The operator $D$ is a modification of the version $\Delta$}. The version $\Delta$} has been introduced in [1][3], and used in [5, (5.6.20)] and [6, (18)], [17], [20, (4)]. The version is symmetric in the following two aspects: for $\rho\in{\cal D}$, ${\Delta}\rho$ is the selfadjoint operator with zero trace that is necessary for the dynamics of $\rho(t)$ in selfadjoint operators with constant trace. The modification $D$ restores the third symmetry with respect to the interchange of $a$ and $a^{\dagger}$.
@@ 4 Injectivity
Let us prove Theorem 2.3 iii). We must show that $\rho=0$ if $D\rho=0$ and $\rho\in{\cal D}$. In the case $D\rho=0$ we have also $\langle\rho,D\rho\rangle=0$. The formula (3.3) gives

### v3

@@ Abstract
We consider dissipation operators used in Quantum Optics for the description of quantum spontaneous emission in the context of damped driven Jaynes–Cummings equations. The equations describe quantised one-mode Maxwell field coupled to a two-level molecule.
We study properties of two basic dissipation operators of Quantum Optics in the framework of the theory of the Hilbert space of the Hilbert-Schmidt operators. We prove that one of the operators is symmetric and nonpositive, while another is neither symmetric nor nonpositive.
@@
Keywords: Quantum Optics; laser; Jaynes–Cummings model; Hamiltonian; density operator; pumping; dissipation; Hilbert space; Hilbert–Schmidt operator; trace.
@@ 1 Jaynes–Cummings equation
The Jaynes–Cummings equation (JCE) is basic model of Quantum Optics. The model without damping and pumping was introduced in 1963 by Jaynes and Cummings [6] (the survey can be found in [9]). The damped driven Jaynes–Cummings equations read as [13, (5.107)]:
@@ 2 Dissipation operators and main results
We will consider the dissipation operators [13, (5.107)] and [4, (5.6.20)]:
@@
$\rm HS$} is the Hilbert space of Hermitian Hilbert–Schmidt operators with the inner product [11, Thm VI.22 (c)]
@@
The operator $D$ is a modification of the version $\Delta$}. The version $\Delta$} has been introduced in [1][3], and used in [4, (5.6.20)] and [5, (18)], [10], [12, (4)]. The version is symmetric in the following two aspects: for $\rho\in{\cal D}$, ${\Delta}\rho$ is the selfadjoint operator with zero trace that is necessary for the dynamics of $\rho(t)$ in selfadjoint operators with constant trace. The modification $D$ restores the third symmetry with respect to the interchange of $a$ and $a^{\dagger}$.
@@ 4 Injectivity
Let us prove Theorem 2.3 iii). We must show that $\rho=0$ if $D\rho=0$ and $\rho\in{\cal D}$. In the case $D\rho=0$, we have also $\langle\rho,D\rho\rangle=0$. The formula (3.3) gives